OpenStudy (anonymous):

A mountain climber hikes 25.5 km from her base camp, going 20.0º north of east. On the second day, she walks 41.0 km in a direction 65.0º north of east, at which point she has reached a summit. Determine the magnitude and direction of her resultant displacement between the base camp and the second camp.

OpenStudy (anonymous):

You need to draw a diagram

OpenStudy (anonymous):

Then connect the end point of the vector values between the starting point and ending point. That is connect from where the climber started to where he finished his trip at the summit.

OpenStudy (anonymous):

Then use the cosine law to calculate the displacement. Sine law may also be used depending on what you like.

OpenStudy (anonymous):

Use the cosine law instead of sine law.

OpenStudy (anonymous):

Perform the following calculation c^2=(25.5)^2+(41)^2+(25.5)(41)cos(115 degrees)= displacement The angle used is calculated from 180-65=115

OpenStudy (anonymous):

let me know if you have any questions